OpenStudy (anonymous):

A load of 5kg stretches a coil spring to a length of 24cm and a load of 8kg stretches to a length of 30cm. find the length of the spring when there is no load

OpenStudy (anonymous):

F= k x F=mg mg = kx

OpenStudy (zzr0ck3r):

are assuming this spring is hanging or on a table?

OpenStudy (anonymous):

physics question but i guess we can help..

OpenStudy (zzr0ck3r):

hook's law is introduced in math as well as physics.

OpenStudy (anonymous):

let "s" be the initial length of the spring x is the total displacement of the spring which means x1 = .24-s and x2 = .30-s

OpenStudy (anonymous):

I don't quite understand but thanks anyways

OpenStudy (anonymous):

5kg stretches a coil spring to a length of 24cm F= mg - equation for weight/force of an object where "m" is mass (kg) and g is the gravitational constant 9.81 m/s^2 F=kx - equation for force of a spring where "k" is the stiffness or spring constant and x is the total displacement of the spring or how far the spring is stretched

OpenStudy (anonymous):

the force of the weight is equal to the force of the spring thus mg = kx

OpenStudy (anonymous):

now applying the first statement 5kg stretches a coil spring to a length of 24cm and substituting hte values in to the equation mg =kx now, from above, x = 24-s substituting all values 5 (9.81) = k (.24 - s)

OpenStudy (anonymous):

now do the same for the 2nd statement

OpenStudy (anonymous):

once you have 2 equations, set both equations equal to k then combine the 2 equations into 1 equation assuming that the stiffness or spring constant of the springs are the same then solve for s which is the initial length of the spring

OpenStudy (anonymous):

ok thanks so much that makes a lot more sense

OpenStudy (anonymous):

no prob